OpenStudy (anonymous):

Twelve people: are given two identical speakers to listen to for differences, if any. Suppose that these people answered by guessing only. Find the probability that three people claim to have heard a difference between the two speakers.

OpenStudy (kropot72):

The probability that a person will claim to have heard a difference is 0.5. The binomial distribution applies, so we have: \[P(3\ out\ of\ 12\ claim\ a\ difference)=12C3\times(0.5)^{12}=you\ can\ calculate\]

OpenStudy (anonymous):

is the ans = 5.3%

OpenStudy (kropot72):

That is not quite the answer that I get.

OpenStudy (kropot72):

12C3 = 220 0.5^12 = 1/4096. So we have: 220/4096 = ?

OpenStudy (anonymous):

0.053

OpenStudy (kropot72):

Not quite. It is 0.0537 which rounds to 0.054.